# VLM-SFD: VLM-Assisted Siamese Flow Diffusion Framework for Dual-Arm Cooperative Manipulation

### v1

@@ V Conclusions
This paper introduces the VLM-Assisted Siamese Flow Diffusion (VLM-SFD) framework for dual-arm cooperative manipulation via imitation learning from only a few human demonstrations. The framework features a Siamese Flow Diffusion Network (SFDNet) that generates two parallel streams of motion flows conditioned on language instructions and visual context, along with a VLM-assisted spatial-temporal task allocation strategy that segments and assigns these motion flows to each robotic arm based on temporal dependencies and spatial constraints. VLM-SFD framework enables robots to learn complex bimanual tasks efficiently and allows direct deployment on real-world robotic hardware without any fine-tuning. Experimental results demonstrate that VLM-SFD outperforms baselines in four challenging real-world tasks.
Limitations VLM-SFD relies on a pre-trained vision-language model for task allocation, which means that relatively heavy computational cost from the VLM can impact performance in complex, dynamic settings. Hence, improving the framework’s robustness to dynamic environment variations and extending it to a wider range of cooperative tasks, and integrating more advanced vision or world models for richer semantic understanding are important directions for future work.
[…]

### v2

@@ V Conclusions
This paper introduces the VLM-Assisted Siamese Flow Diffusion (VLM-SFD) framework for dual-arm cooperative manipulation via imitation learning from only a few human demonstrations. The framework features a Siamese Flow Diffusion Network (SFDNet) that generates two parallel streams of motion flows conditioned on language instructions and visual context, along with a VLM-assisted spatial-temporal task allocation strategy that segments and assigns these motion flows to each manipulator based on temporal dependencies and spatial constraints. VLM-SFD framework empowers to learn complex bimanual tasks efficiently and allows deployment on real-world robotic hardware without fine-tuning. Experimental results demonstrate that VLM-SFD outperforms baselines in four challenging real-world tasks.
Limitations. VLM-SFD relies on a pre-trained vision–language model for task allocation, which introduces non-trivial computational overhead and may affect responsiveness in complex or rapidly changing scenes. The current evaluation is limited to controlled settings and has not fully examined robustness under changes in background, lighting, object types, or occlusion. Real-world execution is also influenced by grasping precision, contact stability, and calibration errors, which can affect performance even with accurate flow predictions. As a learning-based imitation approach, the framework lacks the formal safety guarantees of classical motion planning and prioritizes collision avoidance over tight synchronization in some tasks. Future work will enhance environmental and physical robustness, integrate force/tactile feedback, and combine learning-based planning with classical control for safer, more general dual-arm cooperation.